Mary invested a certain amount of money in a savings account paying 2% simple interest per year when she withdrew her money at the end of 3 years, she received $540 i interest. How much money did Mary place in the savings account?
Answer by fractalier(6550)

The simple interest formula is
I = prt
In this problem we need to find the principal, p, so we have
p = I/rt
Now plug in and get
p = 540/(.02)(3) = $9000